# SynthCoder: A Synthetical Strategy to Tune LLMs for Code Completion

### v2

@@ IV-B1 CrossCodeEval with BM25 Retrieval Strategy
Seed-Coder-8B-Inst 60.2 $\downarrow$8.9% 88.7 $\downarrow$2.8% 84.8 $\downarrow$3.3% 49.1 $\downarrow$5.0% 82.6 $\downarrow$3.0% 78.2 $\downarrow$2.8% 48.6 $\downarrow$1.9% 81.6 $\downarrow$1.5% 76.3 $\downarrow$1.1% 48.5 $\downarrow$6.5% 84.0 $\downarrow$2.8% 79.0 $\downarrow$2.7% 50.7 $\downarrow$5.5% 83.9 $\downarrow$2.5% 79.1 $\downarrow$2.4%
Seed-Coder-8B-Base 61.9 $\downarrow$5.9% 89.1 $\downarrow$2.3% 85.4 $\downarrow$2.5% 52.6 $\uparrow$1.9% 84.5 $\downarrow$0.8% 80.4 -0.0% 49.3 $\downarrow$0.3% 81.9 $\downarrow$1.2% 76.8 $\downarrow$0.4% 50.6 $\downarrow$1.9% 84.8 $\downarrow$1.8% 80.0 $\downarrow$1.4% 52.7 $\downarrow$1.5% 84.7 $\downarrow$1.5% 80.2 $\downarrow$1.1%
SynthCoder-8B-SFT 63.5 $\downarrow$3.3% 89.8 $\downarrow$1.4% 86.3 $\downarrow$1.4% 52.0 $\uparrow$0.8% 84.4 $\downarrow$0.9% 80.2 $\downarrow$0.2% 49.8 $\uparrow$0.6% 82.4 $\downarrow$0.6% 77.4 $\uparrow$0.3% 51.5 -0.0% 85.5 $\downarrow$1.0% 80.8 $\downarrow$0.4% 53.3 $\downarrow$0.4% 85.2 $\downarrow$0.9% 80.7 $\downarrow$0.4%
SynthCoder-8B-DPO-wo-RS 65.4 $\downarrow$0.2% 91.1 -0.0% 87.5 $\downarrow$0.1% 52.9 $\uparrow$2.5% 85.4 $\uparrow$0.3% 80.8 $\uparrow$0.5% 48.7 $\downarrow$1.5% 82.4 $\downarrow$0.6% 76.6 $\downarrow$0.8% 51.7 $\uparrow$0.3% 86.1 $\downarrow$0.3% 80.9 $\downarrow$0.2% 53.6 $\uparrow$0.2% 85.8 $\downarrow$0.2% 80.9 $\downarrow$0.2%
SynthCoder-8B-DPO-RS 65.5 91.1 87.6 51.6 85.1 80.4 49.5 82.9 77.2 51.6 86.3 81.1 53.5 86.0 81.0
IV-B2 CoLT with Dependency-based Strategy Only
The test set of CoLT-132K is a benchmark proposed in the aiXcoder-7B-v2 [10] to evaluate a model’s ability to utilize cross-file context for code completion. It encompasses Python, Java, C/C++, and Go with three completion scenarios: API calls, single-line completions, and span completions, comprising a total of 12,000 samples.
In this work, we use Exact Match (EM) and Edit Similarity (ES) as the evaluation metrics. To specifically assess the model’s ability to leverage dependency-based cross-file context, we exclusively retain the context retrieved through dependency call relationships for each test case. For each model, we employ its corresponding prompt format for completion, with standardized total input context length of 4096 tokens, maximum output length of 200 tokens, and temperature of 0.
As demonstrated in Table IV, SynthCoder-8B shows improvement over its base model in all four programming languages and achieves the best performance among all tested models. This includes outperforming models with larger parameter counts, such as DeepSeek-Coder-V2-Lite-Base, DeepSeek-Coder-V2-Lite-Instruct, and Qwen3-Coder-30B-A3B. The fine-tuned SynthCoder-8B-SFT improves the overall EM and ES by 4.1% and 4.2%, respectively, with gains across all individual metrics, including a maximum improvement of 7.4% on the Go EM score. The aligned SynthCoder-8B-DPO-wo-RS further enhances performance over SynthCoder-8B-SFT on all metrics, with an additional increase of 1.5% in overall EM and 1.4% in overall ES. The final version with repetition suppression, SynthCoder-8B-DPO-RS, achieves the optimal results across all metrics, ultimately boasting a 7.0% improvement in overall EM and a 7.5% improvement in overall ES compared to the base model. It is worth noting that for all metrics on this benchmark, each subsequent stage of our process – fine-tuning, alignment, and repetition suppression – resulted in a consistent improvement in the performance of the SynthCoder-8B model.
@@ IV-B2 CoLT with Dependency-based Strategy Only
Seed-Coder-8B-Inst 42.8 $\downarrow$16.0% 71.7 $\downarrow$12.1% 48.5 $\downarrow$15.2% 76.2 $\downarrow$9.9% 57.7 $\downarrow$8.2% 81.1 $\downarrow$6.3% 46.0 $\downarrow$12.5% 75.1 $\downarrow$8.3% 48.8 $\downarrow$12.7% 76.0 $\downarrow$9.1%
Seed-Coder-8B-Base 46.4 $\downarrow$7.0% 74.2 $\downarrow$8.3% 51.0 $\downarrow$9.7% 76.6 $\downarrow$9.3% 59.6 $\downarrow$4.6% 81.6 $\downarrow$5.7% 48.4 $\downarrow$6.9% 76.3 $\downarrow$6.6% 51.4 $\downarrow$7.0% 77.2 $\downarrow$7.5%
SynthCoder-8B-SFT 47.5 $\downarrow$4.5% 76.6 $\downarrow$4.9% 54.7 $\downarrow$2.3% 81.3 $\downarrow$3.1% 60.8 $\downarrow$2.6% 83.8 $\downarrow$2.9% 50.5 $\downarrow$2.4% 79.4 $\downarrow$2.5% 53.4 $\downarrow$2.9% 80.2 $\downarrow$3.3%
SynthCoder-8B-DPO-wo-RS 48.9 $\downarrow$1.6% 78.6 $\downarrow$2.2% 55.1 $\downarrow$1.6% 82.0 $\downarrow$2.1% 62.1 $\downarrow$0.5% 85.3 $\downarrow$1.1% 50.6 $\downarrow$2.2% 79.6 $\downarrow$2.2% 54.2 $\downarrow$1.4% 81.4 $\downarrow$1.9%
SynthCoder-8B-DPO-RS 49.7 80.3 55.9 83.8 62.4 86.2 51.7 81.4 54.9 82.9
IV-B3 ExecRepoBench
The ExecRepoBench benchmark [^31] https://huggingface.co/datasets/CSJianYang/ExecRepoBench, proposed by the Alibaba team, focuses on evaluating the code completion capabilities of LLMs in real-world software development environments. The dataset is sourced from 50 actively maintained Python repositories on GitHub, comprising a total of 1,200 test cases. For test case construction, the benchmark introduces a multi-level syntax masking method based on Abstract Syntax Trees (AST). It utilizes Tree-sitter to classify syntax nodes into three categories – functions, statements, and expressions – to create test instances. Additionally, it employs non-syntactic construction methods, including snippet, single-line, and multi-line truncation. The benchmark provides not only the intra-file prefix and suffix in a Fill-in-the-Middle (FIM) format but also the relevant cross-file context from the same repository in the form of complete files.
In our reproduction of the ExecRepoEval evaluation, we implemented a rule-based fix to correct indentation errors in the code generated by the models. We use Pass@1 and Edit Similarity (ES) to measure model performance. For each model, we employed its corresponding prompt format for completion, with limits of 4096 tokens for the intra-file context, 4096 tokens for the cross-file context, and a maximum output length of 160 tokens. The temperature was set to 0.
As shown in Table V, SynthCoder-8B achieves a 24.5% improvement in Pass@1 and a 35.3% improvement in ES compared to its base model. Its performance is also significantly higher than that of other 7B and 14B models, including Qwen2.5-Coder-14B-Inst. We did not include Qwen2.5-Coder-Instruct-C, which was proposed in the original ExecRepoBench paper, as its weights have not been publicly released. The fine-tuned SynthCoder-8B-SFT improves upon the base model by 8.4% in Pass@1 and 8.6% in ES. The aligned SynthCoder-8B-DPO-wo-RS further boosts performance by 15.5% in Pass@1 and 25.1% in ES relative to SynthCoder-8B-SFT. The final version with repetition suppression, SynthCoder-8B-DPO-RS, adds an additional 0.6% to Pass@1 and 1.6% to ES, culminating in a total improvement of 24.5% and 35.3% over the original base model.
@@ IV-B3 ExecRepoBench
Seed-Coder-8B-Inst 32.8 $\downarrow$89.3% 38.1 $\downarrow$36.0%
Seed-Coder-8B-Base 45.9 $\downarrow$35.3% 41.6 $\downarrow$24.5%
SynthCoder-8B-SFT 49.0 $\downarrow$26.7% 44.6 $\downarrow$16.1%
SynthCoder-8B-DPO-wo-RS 61.1 $\downarrow$1.6% 51.5 $\downarrow$0.6%
SynthCoder-8B-DPO-RS 62.1 51.8
IV-C Errors of Wrongly Repeating Suffix or Prefix
When applying base models to practical code completion plugins, we observed that models frequently tend to erroneously repeat the preceding (prefix) or succeeding (suffix) context. This behavior can frustrate users and lead them to question the capabilities of both the model and the plugin. To analyze this issue systematically, we focus on a specific scenario: cases where the first non-empty line of the generated content does not exactly match the expected completion, but is identical to the last line of the prefix or the first line of the suffix after whitespace is removed. Our models significantly reduce the probability of such anomalous repetitions. This definition covers instances where the model repeats a context line and then proceeds to generate additional code. Although APIs like OpenAI’s provide various repetition penalty parameters, their effectiveness is limited in the context of Fill-in-the-Middle (FIM) tasks. This is due to the substantially longer input lengths compared to output lengths and the model’s extremely low perplexity when it erroneously repeats the context. To suppress this problem thoroughly, approaches like model-side training or output interception are required.
@@ IV-C Errors of Wrongly Repeating Suffix or Prefix
Seed-Coder-8B-Inst 2.74% $\uparrow$621% 0.43% $\uparrow$377% 2.54% $\uparrow$418%
Seed-Coder-8B-Base 1.67% $\uparrow$339% 2.23% $\uparrow$2377% 1.77% $\uparrow$261%
SynthCoder-SFT 0.60% $\uparrow$57% 2.58% $\uparrow$2766% 1.42% $\uparrow$189%
SynthCoder-DPO-wo-RS 0.52% $\uparrow$36% 0.60% $\uparrow$566% 0.50% $\uparrow$2%
SynthCoder-DPO-RS 0.38% 0.09% 0.49%
@@
Seed-Coder-8B-Inst 0.33% $\uparrow$560% 0.09% -0% 0.08% $\uparrow$300%
Seed-Coder-8B-Base 0.56% $\uparrow$1020% 0.09% -0% 0.08% $\uparrow$300%
SynthCoder-SFT 0.15% $\uparrow$200% 0.34% $\uparrow$278% 0.16% $\uparrow$700%
SynthCoder-DPO-wo-RS 0.10% $\uparrow$100% 0.17% $\uparrow$89% 0.08% $\uparrow$300%
SynthCoder-DPO-RS 0.05% 0.09% 0.02%
@@
We analyzed the frequency of erroneous suffix and prefix repetitions in the generated outputs across the aforementioned datasets, with the results presented in Table VI and Table VII, respectively. The CrossCodeEval dataset is omitted from these tables because its ground-truth completions do not contain full lines, meaning the specific type of whole-line repetition we define does not occur. We observed that for models using either the Prefix-Suffix-Middle (PSM) FIM format (e.g., the Qwen-Coder and DeepSeek series) or the Suffix-Prefix-Middle (SPM) FIM format (e.g., the Seed-Coder series), the frequency of erroneously repeating the suffix is generally higher than that of repeating the prefix. The StarCoder2-7B model exhibits a particularly high repetition rate, reaching a suffix repetition rate of 30.58% on the ExecRepoEval benchmark. While the prefix repetition rate of DeepSeek-Coder is generally lower than that of the Qwen-Coder series, the opposite is true for their suffix repetition rates. Noteworthly, The Seed-Coder series of models demonstrates lower repetition rates compared to the other model families.
We also noted that comparing the model before and after fine-tuning (i.e., Seed-Coder-8B-Base vs. SynthCoder-SFT), the repetition rates exhibit fluctuations of varying magnitudes and directions across different test sets. Interestingly, even without the inclusion of specialized repetition-suppression data, the repetition rate of SynthCoder-DPO-wo-RS is lower than that of SynthCoder-SFT. This is because outputs containing repetitions that are generated and rejected during the sampling process are subsequently used as the negative examples in the preference pairs for DPO training. After incorporating an additional 11% of repetition-suppression data, SynthCoder-DPO-RS demonstrated a significantly greater reduction in this behavior. It achieved a prefix repetition rate below 0.5% and a suffix repetition rate below 0.1% across all datasets, marking the best performance among the evaluated models.
@@ V-A Gap Between Benchmarks and Application
We also observed significant variations in output length among different models given the same input. Ideally, generated code should be concise and to the point, but this is often not the case. Models without fine-tuning are prone to generating excessively long outputs, either by endlessly repeating permutations of the surrounding code or by arbitrarily adding extensive, non-essential docstrings. This can lead to the model reaching its maximum generation length, resulting in unusable code. This behavior is particularly pronounced in the Qwen-Coder series of models and is frequently observed in the ExecRepoEval benchmark. For developers, such unbounded generation not only needlessly consumes computational resources on the model side but also wastes their time and mental energy waiting for and evaluating the results. This underscores the importance of constructing training corpora based on syntax tree nodes and user behavior patterns, which enables the model to generate more targeted and appropriately sized completions. These crucial capabilities are not adequately captured by current evaluation benchmarks.
@@ V-B About the Context Ranges
V-C Next Edit Suggestions
Next Edit Prediction infers a developer’s intent by analyzing the code context and historical edit sequences. It can predict the content and location of the next edit without requiring natural language input, thereby preventing interruptions to the workflow [45]. For instance, during repetitive tasks such as API updates or identifier renaming, it can proactively suggest coherent subsequent edits based on the user’s initial modification. This feature has already been implemented in GitHub Copilot [^32] https://code.visualstudio.com/blogs/2025/02/12/next-edit-suggestions, where it works in synergy with traditional code completion capabilities.
The core of the edit prediction task is to modify existing code: whether to fix bugs or iterate on features, rather than to rewrite it entirely. This area has been a subject of extensive academic research. For example, CoditT5 [46] introduced a pre-trained model specifically designed for code editing. Modit [47] developed a multi-modal editing engine that integrates edit location, code context, and commit information. More recent approaches like jLED [48] employ a joint training methodology, enabling the model to simultaneously learn where to edit and how to edit, providing a more practical solution for real-world editing scenarios. Collectively, these research efforts have advanced the ability of models to follow precise instructions and execute complex code modifications.
In contrast to the fine-grained task of performing insertions, deletions, and modifications at the cursor, many current methods for evaluating the coding capabilities of large models focus on code generation, which does not fully capture the nuances of edit prediction. For example, mainstream function-level benchmarks, including HumanEval [49] and MBPP [50], require models to generate complete code blocks based on function signatures, docstrings, or functional descriptions. The limitation of such benchmarks is that they measure a model’s ability to comprehend high-level instructions and generate code from scratch, rather than its capacity for making precise, intelligent, and incremental modifications within the complex context of existing code. Meanwhile, existing code completion datasets primarily focus on token-level completion at the cursor based on the surrounding code context; they do not incorporate edit history, making it difficult to assess a model’s ability to reason by analogy from past edits. Looking ahead, models such as zeta [^33] https://huggingface.co/zed-industries/zeta have already successfully integrated FIM-based code completion with edit prediction, achieving consistently strong performance across both capabilities. This integration will enable models to become more intelligent and better aligned with developer workflows, driving a paradigm shift in code completion from a reactive response to a proactive prediction model that incorporates edit suggestions.

### v3

@@ IV-B2 CrossCodeEval with BM25 Retrieval Strategy
Seed-Coder-8B-Inst 60.2 $\uparrow$8.9% 88.7 $\uparrow$2.8% 84.8 $\uparrow$3.3% 49.1 $\uparrow$5.0% 82.6 $\uparrow$3.0% 78.2 $\uparrow$2.8% 48.6 $\uparrow$1.9% 81.6 $\uparrow$1.5% 76.3 $\uparrow$1.1% 48.5 $\uparrow$6.5% 84.0 $\uparrow$2.8% 79.0 $\uparrow$2.7% 50.7 $\uparrow$5.5% 83.9 $\uparrow$2.5% 79.1 $\uparrow$2.4%
Seed-Coder-8B-Base 61.9 $\uparrow$5.9% 89.1 $\uparrow$2.3% 85.4 $\uparrow$2.5% 52.6 $\downarrow$1.9% 84.5 $\uparrow$0.8% 80.4 -0.0% 49.3 $\uparrow$0.3% 81.9 $\uparrow$1.2% 76.8 $\uparrow$0.4% 50.6 $\uparrow$1.9% 84.8 $\uparrow$1.8% 80.0 $\uparrow$1.4% 52.7 $\uparrow$1.5% 84.7 $\uparrow$1.5% 80.2 $\uparrow$1.1%
SynthCoder-8B-SFT 63.5 $\uparrow$3.3% 89.8 $\uparrow$1.4% 86.3 $\uparrow$1.4% 52.0 $\downarrow$0.8% 84.4 $\uparrow$0.9% 80.2 $\uparrow$0.2% 49.8 $\downarrow$0.6% 82.4 $\uparrow$0.6% 77.4 $\downarrow$0.3% 51.5 -0.0% 85.5 $\uparrow$1.0% 80.8 $\uparrow$0.4% 53.3 $\uparrow$0.4% 85.2 $\uparrow$0.9% 80.7 $\uparrow$0.4%
SynthCoder-8B-DPO-wo-RS 65.4 $\uparrow$0.2% 91.1 -0.0% 87.5 $\uparrow$0.1% 52.9 $\downarrow$2.5% 85.4 $\downarrow$0.3% 80.8 $\downarrow$0.5% 48.7 $\uparrow$1.5% 82.4 $\uparrow$0.6% 76.6 $\uparrow$0.8% 51.7 $\downarrow$0.3% 86.1 $\uparrow$0.3% 80.9 $\uparrow$0.2% 53.6 $\downarrow$0.2% 85.8 $\uparrow$0.2% 80.9 $\uparrow$0.2%
SynthCoder-8B-DPO 65.5 91.1 87.6 51.6 85.1 80.4 49.5 82.9 77.2 51.6 86.3 81.1 53.5 86.0 81.0
IV-B3 CoLT with Dependency-based Strategy Only
As demonstrated in Table IV, SynthCoder-8B-DPO shows improvement over its base model in all four programming languages and achieves the best performance among all tested models. This includes outperforming models with larger parameter counts, such as DeepSeek-Coder-V2-Lite-Base, DeepSeek-Coder-V2-Lite-Instruct, and Qwen3-Coder-30B-A3B. The fine-tuned SynthCoder-8B-SFT improves the overall EM and ES by 3.9% and 3.9%, respectively, with gains across all individual metrics, including a maximum improvement of 7.3% on the Go EM score. The aligned SynthCoder-8B-DPO-wo-RS further enhances performance over SynthCoder-8B-SFT on all metrics, with an additional increase of 1.5% in overall EM and 1.0% in overall ES. The final version with repetition suppression, SynthCoder-8B-DPO, achieves the optimal results across all metrics, ultimately boasting a 7.0% improvement in overall EM and a 7.5% improvement in overall ES compared to the base model. It is worth noting that for all metrics on this benchmark, each subsequent stage of our process resulted in a consistent improvement in the final.
@@ IV-B3 CoLT with Dependency-based Strategy Only
Seed-Coder-8B-Inst 42.8 $\uparrow$16.0% 71.7 $\uparrow$12.1% 48.5 $\uparrow$15.2% 76.2 $\uparrow$9.9% 57.7 $\uparrow$8.2% 81.1 $\uparrow$6.3% 46.0 $\uparrow$12.5% 75.1 $\uparrow$8.3% 48.8 $\uparrow$12.7% 76.0 $\uparrow$9.1%
Seed-Coder-8B-Base 46.4 $\uparrow$7.0% 74.2 $\uparrow$8.3% 51.0 $\uparrow$9.7% 76.6 $\uparrow$9.3% 59.6 $\uparrow$4.6% 81.6 $\uparrow$5.7% 48.4 $\uparrow$6.9% 76.3 $\uparrow$6.6% 51.4 $\uparrow$7.0% 77.2 $\uparrow$7.5%
SynthCoder-8B-SFT 47.5 $\uparrow$4.5% 76.6 $\uparrow$4.9% 54.7 $\uparrow$2.3% 81.3 $\uparrow$3.1% 60.8 $\uparrow$2.6% 83.8 $\uparrow$2.9% 50.5 $\uparrow$2.4% 79.4 $\uparrow$2.5% 53.4 $\uparrow$2.9% 80.2 $\uparrow$3.3%
SynthCoder-8B-DPO-wo-RS 48.9 $\uparrow$1.6% 78.6 $\uparrow$2.2% 55.1 $\uparrow$1.6% 82.0 $\uparrow$2.1% 62.1 $\uparrow$0.5% 85.3 $\uparrow$1.1% 50.6 $\uparrow$2.2% 79.6 $\uparrow$2.2% 54.2 $\uparrow$1.4% 81.4 $\uparrow$1.9%
SynthCoder-8B-DPO 49.7 80.3 55.9 83.8 62.4 86.2 51.7 81.4 54.9 82.9
IV-B4 ExecRepoBench
As shown in Table V, SynthCoder-8B-DPO achieves significant improvement in Pass@1 and ES compared to its base model, higher than that of other 7B models. We did not include Qwen2.5-Coder-Instruct-C, which was proposed in the original ExecRepoBench paper, as its weights have not been publicly released. The fine-tuned SynthCoder-8B-SFT improves upon the base model by 7.2% in Pass@1 and 6.8% percentage points in ES. The aligned SynthCoder-8B-DPO-wo-RS further boosts performance by 6.9 percentage points in Pass@1 and 12.5 percentage points in ES relative to SynthCoder-8B-SFT. The final version with repetition suppression, SynthCoder-8B-DPO, adds an additional 0.6% to Pass@1 and 1.6% to ES, culminating in a total improvement of 24.5% and 35.3% over Seed-Coder-8B-Base.
@@ IV-B4 ExecRepoBench
Seed-Coder-8B-Inst 32.8 $\uparrow$89.3% 38.1 $\uparrow$36.0%
Seed-Coder-8B-Base 45.9 $\uparrow$35.3% 41.6 $\uparrow$24.5%
SynthCoder-8B-SFT 49.0 $\uparrow$26.7% 44.6 $\uparrow$16.1%
SynthCoder-8B-DPO-wo-RS 61.1 $\uparrow$1.6% 51.5 $\uparrow$0.6%
SynthCoder-8B-DPO 62.1 51.8
IV-C Tendency to Repeat Context
When applying base models to practical code completion plugins, we observed that models frequently tend to erroneously repeat the preceding (prefix) or succeeding (suffix) context. This behavior can frustrate users and lead them to question the capabilities of both the model and the plugin. To analyze this issue systematically, we focus on a specific scenario: cases where the first non-empty line of the generated content does not exactly match the expected completion’s first non-empty line, but is identical to the last non-empty line of the prefix or the first non-empty line of the suffix after whitespace is removed. Our models significantly reduce the probability of such anomalous repetitions. This definition covers instances where the model repeats a context line and then proceeds to generate additional code. Although APIs like OpenAI’s provide various repetition penalty parameters, their effectiveness is limited in the context of Fill-in-the-Middle (FIM) tasks. This is due to the substantially longer input lengths compared to output lengths and the model’s extremely low perplexity when it erroneously repeats the context. To suppress this problem thoroughly, approaches like model-side training or output interception are required.
@@ IV-C Tendency to Repeat Context
Seed-Coder-8B-Inst 2.74% 0.43% 2.54%
Seed-Coder-8B-Base 1.67% 2.23% 1.77%
SynthCoder-SFT 0.60% 2.58% 1.42%
SynthCoder-DPO-wo-RS 0.52% 0.60% 0.50%
SynthCoder-8B-DPO 0.38% 0.09% 0.49%
@@
Seed-Coder-8B-Inst 0.33% 0.09% 0.08%
Seed-Coder-8B-Base 0.56% 0.09% 0.08%
SynthCoder-SFT 0.15% 0.34% 0.16%
SynthCoder-DPO-wo-RS 0.10% 0.17% 0.08%
SynthCoder-8B-DPO 0.05% 0.09% 0.02%
@@
We analyzed the frequency of erroneous suffix and prefix repetitions across the evaluated datasets, with detailed results presented in Table VI and Table VII. The CrossCodeEval dataset was excluded from this analysis because its completion postitions are always in the middle of a line, which precludes the specific type of whole-line repetition under investigation. Particularly pronounced for StarCoder2-7B, which reached a suffix repetition rate of 30.58% on the ExecRepoEval benchmark. Furthermore, while the DeepSeek-Coder series demonstrates a lower prefix repetition rate than the Qwen-Coder series, its suffix repetition rate is conversely higher.
We also noted that comparing the model before and after fine-tuning (i.e., Seed-Coder-8B-Base vs. SynthCoder-SFT), the repetition rates exhibit fluctuations of varying magnitudes and directions across different test sets. Interestingly, even without the inclusion of specialized repetition-suppression data, the repetition rate of SynthCoder-DPO-wo-RS is lower than that of SynthCoder-SFT. This is because outputs containing repetitions that are generated and rejected during the sampling process are subsequently used as the negative examples in the preference pairs for DPO training. After incorporating an additional 11% of repetition-suppression data, SynthCoder-8B-DPO demonstrated a significantly greater reduction in this behavior. It achieved a prefix repetition rate below 0.5% and a suffix repetition rate below 0.1% across all datasets, marking the best performance among the evaluated models.